An electric generating plant produces electric

energy at a rate of 0.2 GW and a voltage of

60968 V.

If it is desired to transmit this power at

1.21936×105 V, what must be the turns ratio

of the transformer used?

015 (part 2 of 2) 10.0 points

What current would be sent out over the

power lines if it were transmitted at 1.21936×

105 V?

Answer in units of A.
Answer #1

Given in the question,

Power, P = 0.2 GW = 0.2*10^9 W = 2*10^8 W

Input potential difference, V1 = 60968 V

Output potential difference, V2 = 1.21936*10^5 V = 121936 V

(1)   The turns ratio of the transformer used n1 / n2 = V2 / V1 = 121936 / 60968  = 2

(2) For the current would be sent out over the power lines if it were transmitted,

V2*I2 = V1*I1

The initial current in the primary coil,

I1 = P / V1 = 2*10^8 / 60968 = 3280.41 A   

Using value of I1 and other values in the above equation,

121936*I2 = 60968*3280.41 => I2 = 1640.21 A